OpenStudy (anonymous):

Parallelogram FGHI on the coordinate plane below represents the drawing of a horse trail through a local park. In order to build a scale model of the trail, the drawing is enlarged on the coordinate plane. If two corners of the trail are at point A (0, 4) and point D (-6, -5), what is another point that could represent a corner of the trail?

OpenStudy (anonymous):

POSSIBLE ANSWERS: A) (12, -5) B) (9, -5) C) (6, -5) D) (3, -5)

OpenStudy (lgbasallote):

@Kreshnik

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i have no idea why he tagged yuu lol

OpenStudy (anonymous):

lol , I think I know why, let me read the question first. :)

OpenStudy (anonymous):

but can you help me ? oh ohk thanks haha

OpenStudy (lgbasallote):

@Kreshnik is good with this stuff :P he's the best i could find hehe

OpenStudy (anonymous):

ok... first we have a model and we have to sketch the new one, with new coordinates. Now first thing to do is we need to know how much bigger will it be... so we need to find distance between IF and DA and we have: \[\LARGE d=\sqrt{(-6+8)^2+(6-3)^2}\] \[\LARGE d=\sqrt{4+9}\] \[\LARGE d=\sqrt{13}\] so distance between points IF is \(\sqrt{13}\) now let's find distance between points DA \[\LARGE d=\sqrt{(0+6)^2+(4+5)^2}\] \[\LARGE d=\sqrt{36+81 }\] \[\LARGE d=\sqrt{117 }\] \[\LARGE d=3\sqrt{13 }\] and we can see that "NEW" (scale or whatever it is, lol) is 3 times bigger... because IF=\(\sqrt{13}\) and DA=\(3\sqrt{13}\) so every "New" side will be 3 times bigger than the old one... I don't know if you've done this (if you've done , it would be better to let me know.. anyway) now let me think about coordinates :)

OpenStudy (anonymous):

wow youre good haha i can never understand this stuff and you just made it a lot easier by explaining it.. (:

OpenStudy (anonymous):

... which corner it is asking lower one or upper one? (sorry I don't understand it well enough - Not American :( )

OpenStudy (anonymous):

i think its asking for the lower one but im not sure either haha its ohk.. i'll figure it out from here(: tahnsk soo much !! (:

OpenStudy (anonymous):

no no ... wait I'll find it, there's no way out (I wanna test myself just to see if I know it lol ).. just give me e sec. ;)

OpenStudy (anonymous):

haha ohk (:

OpenStudy (anonymous):

lol I got it... here it is.. I counted carefully "distance" between similar sides and I counted "points" where NEXT point could be... so here it is next points would be 12,4 and 6,-5 .... and as long as there's no option for point 12,4 (upper one) you have it (lower one) 6,-5 , I also drew a bit on the photo so it could be more clear.... the answer is C 100% sure... here's the photo : (and sorry for taking so long, I made a mistake at principal , and I had to correct it) ... @lgbasallote see my post, does it looks correct to you or not? :)

OpenStudy (anonymous):

and also my web browser just crashed lol :)

OpenStudy (lgbasallote):

lol i tagged you bc i didnt know how to do it (that or i was lazy to read the problem hihi) so you cant confirm it with me :P

OpenStudy (anonymous):

If you were lazy I'm gonn' beat you up :@:@ ... (I was doing a test lol ) hahaha... :)

OpenStudy (lgbasallote):

lol =))) i really didnt know :P

OpenStudy (anonymous):

@Steph_Rawr352 do you understand it?

OpenStudy (anonymous):

yes i got it(: thanks a lot ! youre really good ! thankks !!

OpenStudy (anonymous):

My pleasure xD